# CRVQ: Channel-relaxed Vector Quantization for Extreme Compression of LLMs

### v1

@@ Abstract
Powerful large language models (LLMs) are increasingly expected to be deployed with lower computational costs, enabling their capabilities on resource-constrained devices. Post-training quantization (PTQ) has emerged as a star approach to achieve this ambition, with best methods compressing weights to less than 2 bit on average. In this paper, we propose Channel-Relaxed Vector Quantization (CRVQ), a novel technique that significantly improves the performance of PTQ baselines at the cost of only minimal additional bits. This state-of-the-art extreme compression method achieves its results through two key innovations: (1) carefully selecting and reordering a very small subset of critical weight channels, and (2) leveraging multiple codebooks to relax the constraint of critical channels. With our method, we demonstrate a 38.9% improvement over the current strongest sub-2-bit PTQ baseline, enabling nearer lossless 1-bit compression. Furthermore, our approach offers flexible customization of quantization bit-width and performance, providing a wider range of deployment options for diverse hardware platforms.
@@ 1 Introduction
Transformer-based [35] large language models (LLMs) have made remarkable strides, with open-source models like LLaMA [32, 33], Phi [1] and Qwen [3, 38] achieving exceptional performance across a wide range of benchmarks [18, 9]. However, their massive sizes present significant challenges for deployment and usability. For instance, the widely adopted and powerful LLaMA2-70B model requires over 140GB when loaded in FP16 format, which makes application on most single-GPUs infeasible. This computational and memory overhead becomes even more prohibitive for resource-constrained devices. As the demand for edge-device LLM inference grows, especially on mobile platforms [20, 1], researchers have turned to model compression techniques [16, 30, 37] to enable efficient deployment without compromising performance noticeably.
Post-Training Quantization (PTQ) has emerged as one of the effective methods for compressing LLMs. As a key direction of model quantization, PTQ directly converts pre-trained models to lower-bit representations with relatively few computations [16, 34, 21]. Due to its advantages in both efficiency and effectiveness, PTQ has garnered broader attention in research and practical applications. Recent breakthrough demonstrates that PTQ can compress weights to 2-bit level while astonishingly maintaining lossless performance [34].
Unfortunately, the performance of powerful PTQ methods rapidly deteriorates when compressing weights to extremely low bit-widths. For instance, GPTQ [16], the most popular method for INT4 quantization, completely fails when weights are compressed to 2-bit. Although OmniQuant [30] outperforms GPTQ at the 2-bit level, it still suffers from noticeable performance degradation. A possible reason is that these methods treat each weight as an isolated and uniformly distributed element, independently quantizing them while ignoring the inherent patterns among weights [14, 34]. This oversight results in suboptimal bit utilization, causing these methods to quickly reach their performance limits. Recently, researchers turn to vector quantization (VQ) to address this challenges in PTQ [5, 34, 14]. By treating weights as vectors rather than isolated scalars, VQ-based approaches have shown promising results, see Figure 1. For example, QuIP# [34] emerged as the first PTQ method to achieve near-lossless performance at the 2-bit level. Similarly, AQLM [14] enables seamless adoption of 1-bit settings while surpassing previous strong baselines [40, 21].
@@
Figure 1: Comparison of different quantization methods. Uniform quantization treats each weight to be quantized as a scalar, whereas vector quantization considers weight segments as vectors. Our proposed CRVQ introduces distinct importance across weight channels, with the more critical channels highlighted in dark colors. We use multi-codebook to quantize these critical channels.
@@
Although VQ-based methods have shown promising performance at the 1-bit level, their noticeable performance degradation remains unsatisfactory. Inspired by classical works such as LLM.int8() [10] and SpQR [12], we aim to treat different weight channels with varying importance in VQ. In this paper, we propose Channel-Relaxed Vector Quantization (CRVQ). Traditional VQ generally assume equal importance across all weight channels [34, 14]. In contrast, we argue that even in VQ, a small subset of critical channels can play a pivotal role in maintaining model performance, see the comparison in Figure 1. CRVQ enables these critical channels to break the limits of the current quantization level. This involves two simple yet effective steps: First, we explore useful criteria for evaluating the importance of channels and reorder them, grouping channels of similar importance closer. Then, a few critical channels are selected and quantized with additional codebooks to enhance their representation. Experiments demonstrate that CRVQ significantly improves the performance of 1-bit PTQ with only a negligible cost in bit-width. On the LLaMA2-7B model, CRVQ achieves a 38.9% reduction in perplexity and a 12.9% improvement in zero-shot accuracy compared to strong baselines. In summary, our contributions are 3-fold:
  • We propose a novel and effective method called CRVQ, which leverages critical weight channels to dramatically improve VQ performance in 1-bit PTQ. Remarkably, our method achieves a 39% reduction in perplexity with a cost of only 0.06 bits compared to the previous state-of-the-art method AQLM on LLaMA2-7B model.
@@
  • Extensive experiments demonstrate that our approach performs well in different models of varying sizes, ranging from 125M to 13B, showcasing its effectiveness.
@@ 2 Related Work
Similar to model pruning [15, 31] and knowledge distillation [2, 19], model quantization is a key technique for compressing LLMs and reducing computational overhead. Based on the quantization object, it can be categorized into weight-only quantization and weight-activation quantization. Weight-only quantization (WxA16) focuses on low-bit weight representation. In contrast, weight-activation quantization (WxAy) quantizes both weights and intermediate activations. Model quantization can also be classified into quantization-aware training (QAT) and post-training quantization (PTQ) based on its integrating stage. As this paper primarily focuses on PTQ, the remainder of this section will concentrate on the features and representative methods of QAT and PTQ.
QAT integrates quantization into the intermediate computations of model training, reducing performance degradation caused by low precision and leading superior results. LLM-QAT [26] introduces a few learnable parameters and employs knowledge distillation to transfer abilities from the original model. QLoRA [11] mitigates the adverse effects of quantization by fine-tuning high-precision low-rank adapters. Recently, researchers explore QAT for more effective extreme compression. EfficientQAT [6] adjusts all model through a novel block-wise fine-tuning, achieving outstanding results at the 2-bit quantization level while reducing computational overhead. OneBit [37] designs a new architecture for 1-bit linear layers, incorporating two high-precision adapter vectors and leveraging knowledge distillation to enable effective 1-bit quantization. BinaryMoS [23] further improves on OneBit by replacing the vectors with a mixture-of-experts paradigm. While effective for extreme quantization, QAT still suffers from certain performance limitations and high computational costs.
PTQ transforms LLMs into their low-bit counterparts using solvers and limited calibration data, without requiring extensive training. On the one hand, PTQ has made remarkable progress in extreme compression scenarios. GPTQ [16] iteratively quantizes columns while adjusting other weights to preserve accuracy, achieving 4-bit quantization. QuIP# [34] employs compact E8P codebooks and incoherence processing to enable high-precision 2-bit VQ. Similarly, AQLM [14], a landmark in extreme compression via VQ, leverages learnable codebooks that effectively adapt to the 1-bit setting, leading promising results. Additionally, BiLLM [21] achieves binarization by combining weight selection and residual approximation. Congener works also include SliM-LLM [22], ARB-LLM [24], STB-LLM [13] and Pt-BitNet [17], etc. On the other hand, the idea of mixed precision has been widely applied in quantization research. LLM.int8() [10] is an early approach that proposed mixed-precision decomposition to preserve the functionality of critical channels. Other notable works employing special handling of outlier channels include SmoothQuant [36], AWQ [25] and SpQR [12], etc. Our work pushes the boundaries of 1-bit PTQ by enhancing VQ. Unlike existing methods, we are the first to explore the varying importance of channels within VQ, even though these channels are not adjacent. This novel perspective sets our approach apart from prior work.
@@ 3.1 Vector Quantization
Recently, the application of VQ in extreme PTQ is attractive and is considered as promising approach for pushing the boundaries of PTQ. Unlike uniform quantization, which treats each weight as an independent scalar, VQ partitions the weight matrix into vectors and uses the vector as the basic unit for quantization. Let $\{\mathbf{w}_{i}\}$ represent the set of d-dimensional vectors obtained by partitioning the weight matrix $\mathbf{W}$. VQ applies the K-means algorithm to return the cluster centers of the vectors, which are subsequently used as the codebook:
$\{\mathbf{c}_{i}\}=K\textrm{-means}(\{\mathbf{w}_{j}\}).$ (1)
If each vector in $\{\mathbf{c}_{i}\}$ is encoded using $e$-bit binary number, the codebook $\{\mathbf{c}_{i}\}$ contains $2^{e}$ vectors. Naturally, the capacity of the codebook increases with $e$, leading to better approximation of $\mathbf{W}$. After the codebook is constructed, each vector in $\{\mathbf{w}_{i}\}$ is replaced with its closest vector from $\{\mathbf{c}_{i}\}$. Now, the original $d$-dimensional vector $\mathbf{w}$ is effectively encoded into an $e$-bit binary code.
In classical quantization methods, the quantization process is performed layer by layer. However, this greedy strategy often yields suboptimal results under extreme compression scenarios. To further mitigate the model sensitivity to extreme compression, end-to-end fine-tuning a few parameters is considered highly beneficial [34, 14]. While end-to-end fine-tuning increases the runtime of PTQ, it remains more efficient compared to most QAT methods.
VQ has shown successful applications in extreme compression. QuIP# [34] utilizes a compact, predefined E8P codebook instead of traditional K-means clustering and incorporates incoherence processing to mitigate inter-channel correlations, achieving high-precision 2-bit quantization for the first time. AQLM [14], on the other hand, employs multiple codebooks along with iteratively updated encodings and global fine-tuning, substantially reducing the accuracy loss from extreme compression.
@@ 3.2 From AQLM to 1-bit Quantization
Compared to QuIP#, AQLM are more adaptable for the 1-bit quantization level, as it avoids the challenge of relying on predefined dense codebooks. Let the number of codebooks is $m$, the average bit-width $n$ of a quantized $4096\times 4096$ matrix can be calculated as follows:
@@
1-bit quantized models can be built by selecting different combinations of parameters. Figure 2 illustrates the bit-width contour corresponding to two different codebook number $m$. It can be observed that when the codebook bit-width $e$ exceeds 12 in the left or 10 in the right, the 1-bit contour line rises sharply. It is caused by the excessive codebook storage overhead, reducing the proportion of bit-width allocated to weights. Consequently, larger values of $e$ are avoided.
@@
Figure 2: Visualization of Equation 2 under varying numbers of codebooks $m$. The curves represent bit-width contours based on the vector dimension $d$ and the codebook bit-width $e$. The left (a) and right (b) plots correspond to $m=1$ and $m=2$, respectively.
We adopt a commonly used value of $e=8$, which limits the feasible parameter settings for 1-bit quantization to $m=1,d=8$ or $m=2,d=16$. In this work, we choose the configuration with a smaller bit-width, $m=1,d=8,e=8$. As we will demonstrate later, CRVQ also performs as expected under the $m=2,d=16$ setting.
@@ 4 Approach
This section shows our design for channel-relaxed quantization. We start with a short discussion of how to measure the channel importance in Section 4.1 and then demonstrate how to adapt different importance into VQ in Section 4.2 and Section 4.3.
@@ 4.1 Measuring Channel Importance
It is widely recognized in recent years that different weight channels exhibit varying levels of sensitivity [39, 36, 12]. Consequently, their contributions during the quantization process are not uniform. A considerable number of modern PTQ [12, 10, 36] studies leverage this insight by adopting different treatment of weights, leading to improved results. However, the criteria for measuring weight sensitivity vary across different works.
The Hessian metric serves as a widely used benchmark for evaluating weight sensitivity. Inspired by SpQR[12], we employ an importance criterion that jointly considers quantization error and activation magnitude, with the latter derived from the Hessian proxy. Consequently, the importance $I_{i}$ of a weight channel $\mathbf{W}_{:i}$ can be formulated as:
$I_{i}=\max_{j}\frac{[w_{ji}-\mathrm{VQ}(w_{ji})]^{2}}{2[\mathbf{XX}^{\mathrm{T}}]^{-1}_{ii}},$ (3)
where $\mathbf{X}$ denotes as the activation that multiply weight $\mathbf{W}$ to be quantized. When calculating the importance $I_{i}$, the quantization error is first obtained by pre-quantizing the weights using VQ. Furthermore, some studies [39, 21] also observe that weights exhibit row-wise salients, leading to varying sensitivity to quantization. Intuitively, the presence of salient weights may pose greater challenges for effective quantization. Thus, channel importance can be defined as $I_{i}=\max_{j}w_{ji}^{2}$ as well. The optimal selection in CRVQ will be discussed later. Once the metric $I_{i}$ for each weight channel is computed, the importance of each channel can be determined.

### v2

@@ Abstract
Powerful large language models (LLMs) are increasingly expected to be deployed with lower computational costs, enabling their capabilities on resource-constrained devices. Post-training quantization (PTQ) has emerged as a star approach to achieve this ambition, with best methods compressing weights to less than 2 bit on average. In this paper, we propose Channel-Relaxed Vector Quantization (CRVQ), a novel technique that significantly improves the performance of PTQ baselines at the cost of only minimal additional bits. This state-of-the-art extreme compression method achieves its results through two key innovations: (1) carefully selecting and reordering a very small subset of critical weight channels, and (2) leveraging extended codebooks to relax the constraint of critical channels. With our method, we demonstrate a 38.9% improvement over the current strongest sub-2-bit PTQ baseline, enabling nearer lossless 1-bit compression. Furthermore, our approach offers flexible customization of quantization bit-width and performance, providing a wider range of deployment options for diverse hardware platforms.
CRVQ: Channel-Relaxed Vector Quantization
for Extreme Compression of LLMs
Yuzhuang Xu  Shiyu Ji  Qingfu Zhu  Wanxiang Che🖂 Research Center for Social Computing and Interactive Robotics (SCIR) Harbin Institute of Technology, Harbin, China {xyz, car}@ir.hit.edu.cn
@@ 1 Introduction
Large language models (LLMs) have made remarkable strides, with open-source models like LLaMA (Touvron et al., 2023a, b), Phi (Abdin et al., 2024) and Qwen (Bai et al., 2023; Yang et al., 2024) achieving exceptional performance. However, their massive sizes present challenges for deployment and usability. For instance, the widely adopted and powerful LLaMA2-70B model requires over 140GB when loaded in FP16 format, which makes application on most single-GPUs infeasible. This computational and memory overhead becomes even more prohibitive for resource-constrained devices. As the demand for edge-device LLM inference grows, especially on mobile platforms (Hu et al., 2024; Abdin et al., 2024), researchers have turned to model compression techniques (Frantar et al., 2022; Shao et al., 2024; Xu et al., 2024) to enable efficient deployment without compromising performance noticeably.
(a) uniform quantization
(b) vector quantization
(c) CRVQ
@@
Figure 1: Comparison of different quantization methods. Uniform quantization treats each weight to be quantized as a scalar, whereas vector quantization considers weight segments as vectors. Our proposed CRVQ introduces distinct importance across weight channels, with the more critical channels selected and highlighted in dark colors. We use additional extended codebooks to quantize the vectors formed by these critical channels.
Post-Training Quantization (PTQ) efficiently compresses LLMs by converting pre-trained models to lower-bit formats with relatively few computations (Frantar et al., 2022; Tseng et al., 2024; Huang et al., 2024). Its effectiveness has driven broad research and practical adoption. Notably, recent advances achieve 3-bit compression while astonishingly maintaining lossless performance (Tseng et al., 2024). Unfortunately, the performance of powerful PTQ methods rapidly deteriorates in extremely low bit-widths scenarios. For instance, GPTQ (Frantar et al., 2022), the most popular method for INT4 quantization, completely fails when weights are compressed to 2-bit. A possible reason is that these methods treat each weight as an isolated and uniformly distributed element, independently quantizing them while ignoring the inherent patterns among weights (Egiazarian et al., 2024; Tseng et al., 2024). This oversight results in suboptimal bit utilization, causing these methods to quickly reach their performance limits. Recently, researchers turn to vector quantization (VQ) to address this challenges in PTQ (Chee et al., 2023; Tseng et al., 2024; Egiazarian et al., 2024). By treating weights as vectors rather than isolated scalars (see Figure 1(a) and 1(b)), VQ-based approaches appear to show promising results. For example, AQLM (Egiazarian et al., 2024) enables seamless adoption of 1-bit settings while surpassing previous strong baselines (Yuan et al., 2024; Huang et al., 2024). However, noticeable performance degradation of VQ-based methods remains unsatisfactory at 1-bit level.
@@
Inspired by LLM.int8() (Dettmers et al., 2022) and SpQR (Dettmers et al., 2024), we propose Channel-Relaxed Vector Quantization (CRVQ) in this paper, aiming to treat different weight channels with varying importance in VQ for better quantization. Traditional VQ generally assume equal importance across all weight channels (Tseng et al., 2024; Egiazarian et al., 2024). In contrast, we argue that even in VQ, a small subset of critical channels can play a pivotal role in maintaining model performance. CRVQ enables these critical channels to break the limits of the current quantization level, see the comparison in Figure 1. This involves two simple yet effective steps: First, we explore useful criteria for evaluating the importance of channels and reorder them, grouping channels of similar importance closer. Then, a few critical channels are selected and quantized with additional codebooks to enhance their representation. CRVQ significantly improves 1-bit PTQ, reducing perplexity by 38.9% and boosting zero-shot accuracy by 12.9% on LLaMA2-7B, with only a negligible bit-width cost. In summary, our contributions are 3-fold:
  • We propose CRVQ, a novel and effective 1-bit PTQ method that selects critical weight channels and applies extended codebook fitting. Punching above its weight, it excels in extreme compression.
@@
  • Extensive experiments demonstrate that our approach performs well in different models of varying sizes, ranging from 1.3B to 13B. On LLaMA2-7B, it reduces perplexity by 39% over AQLM with only a 0.06-bit overhead, showcasing its effectiveness.
@@ 2 Related Work
Similar to model pruning (Frantar and Alistarh, 2023; Sun et al., 2024) and knowledge distillation (Agarwal et al., 2024; Hsieh et al., 2023), model quantization is a key technique for compressing LLMs and reducing computational overhead. It can be classified into quantization-aware training (QAT) and post-training quantization (PTQ) based on its integrating stage.
QAT integrates quantization into the intermediate computations of model training, reducing performance degradation caused by low precision and leading superior results (Liu et al., 2024; Dettmers et al., 2023; Chen et al., 2024; Xu et al., 2024; Jo et al., 2024). While effective for extreme quantization, it still suffers from certain performance limitations and high computational costs. As this paper primarily focuses on PTQ, the remainder of this section will concentrate on the features and representative methods of PTQ.
PTQ transforms LLMs into their low-bit counterparts using solvers and limited calibration data, without requiring extensive training. On the one hand, PTQ has made remarkable progress in extreme compression scenarios. GPTQ (Frantar et al., 2022) iteratively quantizes columns while adjusting other weights to preserve accuracy, achieving 4-bit quantization. QuIP# (Tseng et al., 2024) employs compact E8P codebooks and incoherence processing to enable high-precision 2-bit VQ. Similarly, AQLM (Egiazarian et al., 2024) leverages learnable codebooks that can effectively adapt to the 1-bit setting. Additionally, BiLLM (Huang et al., 2024) achieves binarization by combining weight selection and residual approximation. On the other hand, the idea of mixed precision has been widely applied in quantization research. LLM.int8() (Dettmers et al., 2022) is an early approach that proposed mixed-precision decomposition to preserve the functionality of critical channels. Other notable works employing special handling of outlier channels include SmoothQuant (Xiao et al., 2023), AWQ (Lin et al., 2024) and SpQR (Dettmers et al., 2024), etc. Our work pushes the boundaries of 1-bit PTQ by enhancing VQ. Unlike existing methods, we are the first to explore the varying importance of channels within VQ, even though these channels are not adjacent. This novel perspective sets our approach apart from prior work.
@@ 3.1 Vector Quantization
VQ partitions the weight matrix into vectors and uses the vector as the basic quantization unit. Let $\mathcal{S}=\{\mathbf{w}_{i}\in\mathbb{R}^{d}\}$ represent the set of d-dimensional vectors obtained by partitioning $\mathbf{W}^{M\times N}$.
[…]
$\left.j\bmod d=0\right\}.$ (1)
VQ applies the K-means algorithm to return the cluster centers $\mathcal{C}=\{\mathbf{c}_{i}\in\mathbb{R}^{d}\}$, which are subsequently used as the codebook:
$\mathcal{C}=K\textrm{-means}\left(\mathcal{S}\right).$ (2)
If each vector in $\mathcal{C}$ is encoded using $e$-bit binary number, the codebook $\mathcal{C}$ contains $2^{e}$ vectors, i.e. $|\mathcal{C}|=2^{e}$. After the codebook is built, each vector in $\mathcal{S}$ is replaced with its closest vector from $\mathcal{C}$.
$\mathcal{S}\approx\left\{\hat{\mathbf{w}}_{i}=\arg\min_{\mathbf{c}_{j}\in\mathcal{C}}\|\mathbf{w}_{i}-\mathbf{c}_{j}\|_{2}\right\}.$ (3)
Now, the original $d$-dimensional vector $\mathbf{w}$ is encoded into an $e$-bit binary code $q$ as follows:
$q_{i}=\arg\min_{j}\|\mathbf{w}_{i}-\mathbf{c}_{j}\|_{2},\quad q_{i}\in{0,1,\dots,2^{e}-1}.$ (4)
Naturally, the capacity of the codebook increases with $e$, leading to better approximation of $\mathbf{W}$. To reduce approximation error, we can iteratively refine the process above by quantizing the residual $\Delta\mathbf{W}=\mathbf{W}-\hat{\mathbf{W}}$, introducing more codebooks and codes. Beam search is typically used for the optimal multi-step encoding (Egiazarian et al., 2024).
To further mitigate the model sensitivity to extreme compression, block or end-to-end fine-tuning a few parameters is considered highly beneficial (Tseng et al., 2024; Egiazarian et al., 2024).
$\mathcal{L}=\|\mathrm{block}(\mathbf{X})-\mathrm{quantized\_block}(\mathbf{X})\|_{2}.$ (5)
Let $\mathcal{L}$ be the loss and $\mathbf{X}$ the block input here. While end-to-end fine-tuning slows PTQ, it remains more efficient than most QAT methods.
@@ 3.2 From AQLM to 1-bit Quantization
Compared to QuIP#, AQLM are more adaptable for the 1-bit quantization level, as it avoids the challenge of relying on predefined dense codebooks. Let the number of codebooks is $m$, the average bit-width $n$ of a quantized $4096\times 4096$ matrix can be calculated as follows (see Section A.1 for details):
@@
1-bit quantized models can be built by selecting different combinations of hyper parameters. Figure 2 illustrates the bit-width curve for various $e,d$ corresponding to two different codebook number $m$, respectively. It can be observed that when the codebook bit-width $e$ exceeds 14, the average bit-width rises sharply. It is caused by the excessive codebook storage overhead, reducing the proportion of bit-width allocated to weights. Consequently, larger values of $e$ are avoided.
@@
Figure 2: Visualization of Equation 6. The curves represent bit-width based on the vector dimension $d$ and the codebook bit-width $e$. The top (a) and bottom (b) plots correspond to $m=1$ and $m=2$, respectively.
We adopt a commonly used value of $e=8$, which limits the feasible settings for 1-bit quantization to $m=1,d=8$ or $m=2,d=16$. In this work, we choose the configuration with a smaller bit-width, $m=1,d=8,e=8$. As we will demonstrate in Section 6.4, CRVQ also performs as expected under the $m=2,d=16$ setting.
@@ 4 Approach
This section shows our design for channel-relaxed quantization. We start with a discussion of how to measure the channel importance in Section 4.1 and then demonstrate how to adapt different importance into VQ in Section 4.2 and 4.3. Finally, we demonstrate the overall algorithm in Section 4.4.
@@ 4.1 Measuring Channel Importance
It is widely recognized in recent years that different weight channels exhibit varying levels of sensitivity (Yao et al., 2022; Xiao et al., 2023; Dettmers et al., 2024). Consequently, the contributions of these channels in the quantization process are not uniform. A considerable number of modern PTQ studies (Dettmers et al., 2024, 2022; Xiao et al., 2023) leverage this insight by adopting different treatment of weights, leading to improved results. However, the criteria for measuring weight sensitivity vary across different works.
The Hessian metric serves as a widely used benchmark for evaluating weight sensitivity. Inspired by SpQR(Dettmers et al., 2024), we first employ an importance criterion that jointly considers quantization error and activation magnitude, with the latter derived from the Hessian proxy. Consequently, the importance $I_{i}$ of a weight channel $\mathbf{W}_{:,i}$ can be formulated as:
$I_{i}=\max_{j}\frac{\bigl{[}w_{ji}-\mathrm{VQ}\left(w_{ji}\right)\bigr{]}^{2}}{2\left[\mathbf{XX}^{\mathrm{T}}\right]^{-1}_{ii}},$ (7)
where $\mathbf{X}$ denotes as the activation that multiply weight $\mathbf{W}$ to be quantized, see Section A.2 for details. When calculating the importance $I_{i}$, the quantization error is first obtained by pre-quantizing the weights using VQ and then multiplying activation.
Furthermore, some studies (Yao et al., 2022; Huang et al., 2024) also observe that weights exhibit row-wise salients, leading to varying sensitivity to quantization. Intuitively, the presence of salient weights may pose greater challenges for effective quantization. Thus, channel importance may be defined as $I_{i}=\max_{j}w_{ji}^{2}$ as well.
The optimal selection in CRVQ will be discussed in Section 6.3. Once the metric $I_{i}$ for each weight channel is computed, the importance of each channel can be determined.